OpenStudy (anonymous):

The slope of the tangent line to the curve f(x)=x^3-5 at x=3 is 22 true or false? Steps to solve this please

OpenStudy (anonymous):

false

OpenStudy (anonymous):

have u learned the derivative rules?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Differentiate the function: \[f(x) = x^3 - 5\\ f'(x) = 3x^2\] Now, find the slope of the tangent line: \[f'(3) = 3(3)^2 = 3^3 = 27\]

OpenStudy (anonymous):

ok so u kno the derivative is 3x^2 so plug in 3 and youll end up with 18

OpenStudy (anonymous):

whoops 27

OpenStudy (anonymous):

okay thanks guys!